OpenStudy (anonymous):

A toy manufacturer produces the same toy in 4 different colors: red, blue, green, and yellow. The toy manufacturer has 5 production lines dedicated to the red version, 5 production lines dedicated to the blue version, 4 production lines dedicated to the green version, and 2 production lines dedicated to the yellow version. A toy has an equal chance of traveling down any production line. What is the probability that the toy will not be green?

OpenStudy (anonymous):

i have read this not three times and i have no idea what it means but i am going to make a guess

OpenStudy (anonymous):

\[5+5+4+2=16\] productions lines, and of those \(12\) are not green my best guess is therefore \[\frac{12}{16}=\frac{3}{4}\] but honestly the line "A toy has an equal chance of traveling down any production line. " makes no sense to me